OpenStudy (anonymous):

Let c be any constant number. Which of the following will always be perpendicular to −3x + y = 2?

OpenStudy (anonymous):

y=2-3x

OpenStudy (anonymous):

Alternate forms is y=2-3x 3x+y-2=0

OpenStudy (anonymous):

It's one of these: y = 3x + c y = x + c y = −3x + c y = x + c

OpenStudy (anonymous):

I would have to say the first one

OpenStudy (bahrom7893):

-3x+c

OpenStudy (bahrom7893):

oh wait NO. kumar's right

OpenStudy (bahrom7893):

I thought it said parallel

OpenStudy (anonymous):

Oh, okay.. Thanks a bunch! :DD

OpenStudy (anonymous):

\[−3x + y = 2\rightarrow m=3\]So the slope of any perpendicular line is \[m _{\perp}=-1/3\]To get a function in terms of c as requested in the question, use general form:\[x+3y=c,c \in \mathbb{R}\]

OpenStudy (radar):

-3x+y=2 y=3x+2, note that slope is 3, a perpendicular line to this will have a negative reciprocal or a slope of -1/3

OpenStudy (radar):

I think you need more choices ??? I don't see a slope meeting the requirements!

OpenStudy (anonymous):

None of your multiple choices are perpendicular to the original line.

OpenStudy (anonymous):

... Those are the only four that were given. ;-;

OpenStudy (radar):

You sure the third one down, doesn't have a 1 as a numerator and a -3 as the denominator like (-1/3)x

OpenStudy (anonymous):

Oh. They're actually: y = 3x + c y = 1/3x + c y = −3x + c y = -1/3x + c

OpenStudy (anonymous):

Well, which one has a slope of -1/3?

OpenStudy (anonymous):

... The last one? xD